# The dimension and Bose distance of some BCH codes of length $\frac{q^{m}-1}{\lambda}$

### v1

@@ Abstract
BCH codes are important error correction codes, widely utilized due to their robust algebraic structure, multi-error correcting capability, and efficient decoding algorithms. Despite their practical importance and extensive study, their parameters, including dimension, minimum distance and Bose distance, remain largely unknown in general. This paper addresses this challenge by investigating the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$ over the finite field $\mathbb{F}_{q}$, where $\lambda$ is a positive divisor of q1q-1. Specifically, for narrow-sense BCH codes of this length with $m\geq 4$, we derive explicit formulas for their dimension for designed distance $2\leq\delta\leq(q^{\lfloor(2m-1)/3\rfloor+1}-1)/{\lambda}+1$. We also provide explicit formulas for their Bose distance in the range $2\leq\delta\leq(q^{\lfloor(2m-1)/3\rfloor+1}-1)/{\lambda}$. These ranges for $\delta$ are notably larger than the previously known results for this class of BCH codes. Furthermore, we extend these findings to determine the dimension and Bose distance for certain non-narrow-sense BCH codes of the same length. Applying our results, we identify several BCH codes with good parameters.
October 2, 2025
@@ I Introduction
and each code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ with a subset of the quotient ring $\mathbb{F}_{q}[x]/(x^{n}-1).$ In this way, a code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is a cyclic code if and only if it is an ideal of the quotient ring $\mathbb{F}_{q}[x]/(x^{n}-1).$ Note that each ideal of $\mathbb{F}_{q}[x]/(x^{n}-1)$ is principal. Therefore, a cyclic code $\mathcal{C}\subseteq\mathbb{F}_{q}[x]/(x^{n}-1)$ can be generated by a monic polynomial $g(x)$, denoted as $\mathcal{C}=\left\langle g(x)\right\rangle.$ Moreover, the polynomial $g(x)$ is a divisor of $x^{n}-1$. The generator $g(x)$ is called the generator polynomial of $\mathcal{C}$, and $h(x)=(x^{n}-1)/{g(x)}$ is called the parity-check polynomial of $\mathcal{C}$.
Suppose that $n$ is an integer such that $\mathrm{gcd}(n,q)=1$. Denote $m=\mathrm{ord}_{n}(q)$, i.e., the smallest integer such that $q^{m}\equiv 1\ (\mathrm{mod}\ n).$ Let $\alpha$ be a primitive element of $\mathbb{F}_{q^{m}}$. Then $\beta=\alpha^{\frac{q^{m}-1}{n}}$ is a primitive $n$-th root of unity. This leads to the factorization $x^{n}-1=\prod\limits_{i=0}^{n-1}(x-\beta^{i}).$ For each integer $i\in[0,n-1]$, we denote by $m_{i}(x)$ the minimal polynomial of $\beta^{i}$ over $\mathbb{F}_{q}$. A cyclic code of length $n$ over $\mathbb{F}_{q}$ is called a BCH code with designed distance $\delta$ if its generator polynomial takes the form
$\mathrm{lcm}(m_{b}(x),m_{b+1}(x),\ldots,m_{b+\delta-2}(x))$
for some integers $b$ and $2\leq\delta\leq n$, where $\mathrm{lcm}$ denotes the least common multiple of the polynomials. We denote by $\mathcal{C}_{(q,n,\delta,b)}$ such a BCH code. If $b=1$, it is called a narrow-sense BCH code, simply denoted by $\mathcal{C}_{(q,n,\delta)}$. If $n=q^{m}-1$, then it is called a primitive BCH code. Note that $\mathcal{C}_{(q,n,\delta,b)}$ and $\mathcal{C}_{(q,n,\delta,^{{}^{\prime}}b)}$ may be the identical for distinct $\delta$ and $\delta^{{}^{\prime}}$. The Bose distance of $\mathcal{C}_{(q,n,\delta,b)}$, denoted by $d_{B}$ or $d_{B}(\mathcal{C}_{(q,n,\delta,b)})$, is the largest integer such that $\mathcal{C}_{(q,n,\delta,b)}=\mathcal{C}_{(q,n,d_{B},b)}$.
@@
This paper is dedicated to the investigation of the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$, denoted by $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta,b)}$, where $\lambda$ is a positive divisor of q1q-1. This class of BCH codes includes primitive BCH codes when $\lambda=1$. To date, these fundamental parameters are precisely known only for limited cases. Most existing results focus on the case where $\lambda=1$ and $b=1$, which corresponds to narrow-sense primitive BCH codes. For a comprehensive overview of the parameters of such codes, readers are referred to [39, 27, 14, 28, 7, 8]. In contrast, for cases when $\lambda\neq 1$, the understanding is much more limited. Even for certain specific cases, such as $\lambda=2$ [24, 36, 40] and $\lambda=q-1$ [25, 22], the dimension and Bose distance are known for only a few designed distances. For general divisors $\lambda$ of q1q-1, Zhu et al. [40] determined the dimension of narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for designed distances $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$. Recently, Sun [30] determined the dimension and minimum distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for other specific designed distances. Readers may refer to [10] for an excellent survey on known results regarding the parameters of BCH codes.
@@
To extend these results to BCH codes of length $\frac{q^{m}-1}{\lambda}$, we need to investigate $q$-cyclotomic cosets modulo $\frac{q^{m}-1}{\lambda}$. A key and useful observation in this regard is that an integer $a$ is a coset leader modulo $\frac{q^{m}-1}{\lambda}$ if and only if $\lambda a$ is a coset leader modulo $q^{m}-1$, and the size of the $q$-cyclotomic coset modulo $\frac{q^{m}-1}{\lambda}$ of $a$ is equal to the size of the $q$-cyclotomic coset modulo $q^{m}-1$ of $\lambda a$. Consequently, the problem of finding coset leaders and sizes of $q$-cyclotomic coset modulo $\frac{q^{m}-1}{\lambda}$ can be reduced to identifying integers divisible by $\lambda$ that are coset leaders modulo $q^{m}-1$ and determining the sizes of their corresponding $q$-cyclotomic coset modulo $q^{m}-1$.
Building upon this crucial observation and our prior analysis of $q$-cyclotomic cosets modulo $q^{m}-1$, we successfully generalize the results on primitive BCH codes in [39] to BCH codes of length $\frac{q^{m}-1}{\lambda}$. Specifically, for any positive divisor $\lambda$ of q1q-1 and positive integer $m\geq 4$, this paper determines:
@@
It is important to note that the existing knowledge of the dimension for narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ only covers designed distances $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$ and some specific cases.
Our results significantly extend this range, as evidenced by the inequality $\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}\geq\frac{q^{\lfloor(m+1)/2\rfloor}-1}{\lambda}\cdot q^{\lceil(m-4)/6\rceil}$. This implies that the range of $\delta$ for which we provide the dimension of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ is substantially larger than previously established. Additionally, we extend these results to some non-narrow-sense BCH codes of length $\frac{q^{m}-1}{\lambda}$. As illustrations of our main results, we also provide some explicit formulas determining the dimension and Bose distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $\lambda\delta=aq^{h+k}+b$, where $h=\lfloor m/2\rfloor$, $m-2h\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{m-h-k}$, and $q\nmid b$.
This paper is organized as follows. In Sections II and III, we provide essential preliminaries and review our previous results concerning $q$-cyclotomic cosets modulo $q^{m}-1$. Section IV presents several auxiliary lemmas that will be employed in subsequent sections. Based on the theoretical foundations established in Sections II through IV, we then determine the dimension of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$ in Section V and the Bose distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta,b)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ in Section VI, both by providing explicit formulas. Utilizing these formulas, we present some examples of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ and compare them with the tables of the best known linear codes maintained by Markus Grassl at http://www.codetables.de, which is called Database later in this paper. Furthermore, as an illustration of our main results, Section VII applies these formulas to compute the dimension and Bose distance of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ specifically when $\lambda\delta$ takes the form $aq^{h+k}+b$. Following this, Section VIII extends our analysis to determine the dimension and Bose distance of certain non-narrow-sense BCH codes. Moreover, we identify some non-narrow-sense BCH codes that possess optimal parameters. Finally, Section IX concludes the paper.
@@ II Preliminaries
where $\ell_{a}$ is the smallest positive integer such that $aq^{\ell_{a}}\equiv a\pmod{n}$. It is clear that $|C_{n}(a)|=\ell_{a}$, where $|\cdot|$ denotes the size of a set. Moreover, $|C_{a}|$ is a divisor of $m=\mathrm{ord}_{n}(q)$. The smallest integer in the coset $C_{n}(a)$ is called the coset leader of $C_{n}(a)$. For convenience, we occasionally refer to it simply as a coset leader modulo $n$. Let $\ell$ be a positive integer, and let $c,d\in[0,n-1]$ be two integers such that $c\leq d$. We define
@@
In particular, when (c,d)=(0,n1)(c,d)=(0,n-1), we simply denote $\mathcal{L}_{\ell}(c,d)$ by $\mathcal{L}_{\ell}$.
By [17, Theorem 4.1.1], the minimal polynomial $m_{a}(x)$ of $\beta^{a}$ over $\mathbb{F}_{q}$ can be given by
$m_{a}(x)=\prod\limits_{i\in C_{n}(a)}(x-\beta^{i}).$
Hence, the generator polynomial $g(x)$ of the narrow-sense BCH code $\mathcal{C}_{(q,n,\delta)}$ can be expressed as
$g(x)=\prod\limits_{i\in\mathcal{G}}(x-\beta^{i}),\quad\mathcal{G}=\bigcup\limits_{a=1}^{\delta-1}C_{n}(a).$
Consequently, the dimension of $\mathcal{C}_{(q,n,\delta)}$ can be given by
@@
Thus, determining the coset leaders modulo $n$ within $1,\delta-1]$ and the sizes of corresponding cyclotomic cosets allows us to compute the dimension of $\mathcal{C}_{(q,n,\delta)}$. Additionally, for $\delta^{{}^{\prime}}>\delta$, one can easily verify that
$\bigcup\limits_{a=1}^{\delta-1}C_{n}(a)=\bigcup\limits_{a=1}^{\delta^{\prime}-1}C_{n}(a)$
if and only if all the integers in $\left[\delta,\delta^{{}^{\prime}}-1\right]$ are not coset leaders. Therefore, the Bose distance of $\mathcal{C}_{(q,n,\delta)}$ is equal to the smallest coset leader within the range $\delta,n-1]$.
We have a useful observation regarding $q$-cyclotomic cosets, as presented in the following lemma. This result originated from the proof of [40, Lemma 6]. For completeness, we include a proof below.
@@
Noticing that $a\bmod n=a-n\cdot\lfloor\frac{a}{n}\rfloor$ and $\lambda a\bmod\lambda n=\lambda a-\lambda n\cdot\lfloor\frac{\lambda a}{\lambda n}\rfloor$, we can assert that
$\lambda\cdot(a\bmod n)=\lambda a-\lambda n\cdot\lfloor\frac{a}{n}\rfloor=\lambda a-\lambda n\cdot\lfloor\frac{\lambda a}{\lambda n}\rfloor=\lambda a\bmod\lambda n.$
@@
By definition, this holds if and only if $\lambda a$ is the coset leader of $C_{\lambda n}(\lambda a)$. Therefore, the first statment follows.
@@
Let $\mathbb{Z}$ be the set of all integers. For each integer $\lambda$, we denote by $\mathcal{D}_{\lambda}$ the set of integers that are divisible by $\lambda$, that is,
@@
Then one can directly derive the following corollary from the above lemma.

### v2

@@ Abstract
BCH codes are important error correction codes, widely utilized due to their robust algebraic structure, multi-error correcting capability, and efficient decoding algorithms. Despite their practical importance and extensive study, their parameters, including dimension, minimum distance and Bose distance, remain largely unknown in general. This paper addresses this challenge by investigating the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$ over the finite field $\mathbb{F}_{q}$, where $\lambda$ is a positive divisor of q1q-1. Specifically, for narrow-sense BCH codes of this length with $m\geq 4$, we derive explicit formulas for their dimension for designed distance $2\leq\delta\leq(q^{\lfloor(2m-1)/3\rfloor+1}-1)/{\lambda}+1$. We also provide explicit formulas for their Bose distance in the range $2\leq\delta\leq(q^{\lfloor(2m-1)/3\rfloor+1}-1)/{\lambda}$. These ranges for $\delta$ are notably larger than the previously known results for this class of BCH codes. Furthermore, we extend these findings to determine the dimension and Bose distance for certain non-narrow-sense BCH codes of the same length. Several optimal linear codes can be obtained from these BCH codes.
March 10, 2026
@@ I Introduction
and each code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ with a subset of the quotient ring $\mathbb{F}_{q}[x]/(x^{n}-1).$ In this way, a code $\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$ is a cyclic code if and only if it is an ideal of the quotient ring $\mathbb{F}_{q}[x]/(x^{n}-1).$ Note that each ideal of $\mathbb{F}_{q}[x]/(x^{n}-1)$ is principal. Therefore, a cyclic code $\mathcal{C}\subseteq\mathbb{F}_{q}[x]/(x^{n}-1)$ can be generated by a monic polynomial $g(x)$, denoted as $\mathcal{C}=\left\langle g(x)\right\rangle.$ Moreover, the polynomial $g(x)$ is a divisor of $x^{n}-1$. The polynomial $g(x)$ is called the generator polynomial of $\mathcal{C}$, and $h(x)=(x^{n}-1)/{g(x)}$ is called the parity-check polynomial of $\mathcal{C}$.
Suppose that $n$ is a positive integer such that $\mathrm{gcd}(n,q)=1$. Let $m=\mathrm{ord}_{n}(q)$, i.e., the smallest integer such that $q^{m}\equiv 1\ (\mathrm{mod}\ n).$ Let $\alpha$ be a primitive element of the field $\mathbb{F}_{q^{m}}$. Then $\beta=\alpha^{\frac{q^{m}-1}{n}}$ is a primitive $n$-th root of unity. This leads to the factorization $x^{n}-1=\prod\limits_{i=0}^{n-1}(x-\beta^{i}).$ For each integer $i\in[0,n-1]$, we denote by $m_{i}(x)$ the minimal polynomial of $\beta^{i}$ over $\mathbb{F}_{q}$. A cyclic code of length $n$ over $\mathbb{F}_{q}$ is called a BCH code with designed distance $\delta$ if its generator polynomial takes the form
$\mathrm{lcm}(m_{b}(x),m_{b+1}(x),\ldots,m_{b+\delta-2}(x)),$
where $b$ and $\delta$ are integers with $2\leq\delta\leq n$, and $\mathrm{lcm}$ denotes the least common multiple of the polynomials. We denote by $\mathcal{C}_{(q,n,\delta,b)}$ such a BCH code. If $b=1$, it is called a narrow-sense BCH code, simply denoted by $\mathcal{C}_{(q,n,\delta)}$. If $n=q^{m}-1$, then it is called a primitive BCH code. Note that $\mathcal{C}_{(q,n,\delta,b)}$ and $\mathcal{C}_{(q,n,\delta^{{}^{\prime}},b)}$ may be identical even for distinct $\delta$ and $\delta^{{}^{\prime}}$. The Bose distance of $\mathcal{C}_{(q,n,\delta,b)}$, denoted by $d_{B}$ or $d_{B}(\mathcal{C}_{(q,n,\delta,b)})$, is the largest integer such that $\mathcal{C}_{(q,n,\delta,b)}=\mathcal{C}_{(q,n,d_{B},b)}$.
@@
This paper is dedicated to the investigation of the dimension and Bose distance of BCH codes of length $(q^{m}-1)/\lambda$, where $\lambda$ is a positive divisor of q1q-1. This class of BCH codes includes primitive BCH codes when $\lambda=1$. To date, these fundamental parameters are precisely known only for limited cases. Most existing results focus on the case where $\lambda=1$ and $b=1$, which corresponds to narrow-sense primitive BCH codes. For a comprehensive overview of the parameters of such codes, readers are referred to [40, 26, 28, 14, 29, 7, 8].
In contrast, for cases when $\lambda\neq 1$, the understanding is much more limited. Even for some specific cases, such as $\lambda=2$ and $\lambda=q-1$, the dimension and Bose distance are known only for a few designed distances. For general positive divisors $\lambda$ of q1q-1, Zhu et al. [41] determined the dimension of narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for designed distances $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$. Recently, Sun [31] determined the dimension and minimum distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for other specific designed distances. For convenience, we summarize in Table I some known results on the parameters of narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for divisors $\lambda$ of q1q-1 with $\lambda\neq 1$. A check mark ✓indicates that the corresponding parameter is known, while a blank entry indicates that it is unknown. For non-narrow-sense BCH codes of length $(q^{m}-1)/\lambda$, the available results in the literature are even more scarce; see [26] for some related work. Readers may also refer to [10] for an excellent survey on known results regarding the parameters of BCH codes.
Table I: Known results on the parameters of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for divisors $\lambda$ of q1q-1 with $\lambda\neq 1$
$q$ $\lambda$ $m$ $\delta$ $k$ $_{B} dd$ Reference
prime power $\lambda=q-1$ $m\geq 4$ is even $\leq\delta\leq q^{{m}/{2}} [22$]
prime power $\lambda=q-1$ $m\geq 5$ is odd $\delta=aq^{(m-1)/2}+1$ with $1\leq a\leq q-1$ [26]
prime power $\lambda=q-1$ $m\geq 1$ $\delta=2$ [24]
$q\geq 3$ $\lambda=q-1$ $m\geq 1$ $\delta=3$ partially solved [24]
$q=3$ $\lambda=2$ $m\geq 3$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lfloor(m-3)/2\rfloor+i}+1}{2}$, $i=1,2$ [24]
$q$ is odd $\lambda=2$ $m\geq 3$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lfloor(m-3)/2\rfloor+i}+1}{2}$, $i=1,2$ [41]
$q$ is odd $\lambda=2$ $m\geq 2$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lfloor(m-3)/2\rfloor+i}+1}{2}$, $1\leq i\leq\lfloor\frac{(m+11)}{6}\rfloor$ [25]
$q$ is odd $\lambda=2$ $m\geq 4$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lceil\frac{3m}{4}\rceil-1}+q^{\lceil\frac{m}{2}\rceil-1}+q^{\lceil\frac{m}{4}\rceil-1}+1}{2}$ [37]
$q$ is odd $\lambda=2$ $m\geq 7$, $m\equiv 3(\bmod 4)$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\frac{3m-1}{4}}+q^{\frac{m+1}{2}}+q^{\frac{m-11}{4}+t}+1}{2}$, $t=2,3,4$ [37]
$q$ is odd $\lambda=2$ $m\geq 7$, $m\not\equiv 3(\bmod 4)$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lceil\frac{3m}{4}\rceil}+1}{2}$ [37]
$q$ is odd $\lambda=2$ $m=2$ $\delta=\frac{i_{1}q+i_{0}}{2}$, $0\leq i_{1}\leq q-2$, $\max\{i_{0},1\}\leq i_{0}\leq q-1$ and $i_{0}\equiv i_{1}(\bmod 2)$ [37]
$q$ is odd $\lambda=2$ $m=3$ $\delta=\frac{i_{2}q^{2}+i_{1}q+i_{0}}{2}$, $0\leq i_{2}\leq q-2$, $i_{2}\leq i_{1}\leq q-1$, $\max\{i_{2},1\}\leq i_{0}\leq q-1$, $i_{1}+i_{0}\equiv i_{2}(\bmod 2)$ [37]
$q$ is odd $\lambda=2$ $m\geq 4$ $\delta=\frac{q^{m}-q^{m-1}}{2}-\frac{q^{\lfloor(m-3)/2\rfloor+i}+1}{2},1\leq i\leq\lfloor\frac{m+6}{4}\rfloor$ [37]
prime power $\lambda\mid q-1$ $m\geq 3$ is odd $2\leq\delta\leq\frac{q^{(m+1)/2}-1}{\lambda}+1$ [41]
prime power $\lambda\mid q-1$ $m\geq 4$ is even $2\leq\delta\leq\frac{q^{m/2+1}-1}{\lambda}+1$ [41]
prime power $\lambda\mid q-1$ $m\geq 2$ $\delta=\frac{(q-\lambda\ell_{0})q^{m-1-\ell_{1}}-1}{\lambda}$, $0\leq\ell_{1}\leq m-1$ and $0\leq\ell_{0}<\frac{q-1}{\lambda}$ [31]
@@
To extend these results to BCH codes of length $(q^{m}-1)/\lambda$, we need to investigate $q$-cyclotomic cosets modulo $(q^{m}-1)/\lambda$. A key and useful observation in this regard is that an integer $a$ is a coset leader modulo $(q^{m}-1)/\lambda$ if and only if $\lambda a$ is a coset leader modulo $q^{m}-1$, and the size of the $q$-cyclotomic coset modulo $(q^{m}-1)/\lambda$ of $a$ is equal to the size of the $q$-cyclotomic coset modulo $q^{m}-1$ of $\lambda a$. Consequently, the problem of finding coset leaders and sizes of $q$-cyclotomic coset modulo $(q^{m}-1)/\lambda$ can be reduced to identifying integers divisible by $\lambda$ that are coset leaders modulo $q^{m}-1$ and determining the sizes of their corresponding $q$-cyclotomic coset modulo $q^{m}-1$.
Building upon this crucial observation and our prior analysis of $q$-cyclotomic cosets modulo $q^{m}-1$, we successfully generalize our results on primitive BCH codes in [40] to BCH codes of length $(q^{m}-1)/\lambda$. Specifically, for any positive divisor $\lambda$ of q1q-1 and positive integer $m\geq 4$, this paper determines:
@@
It is important to note that the existing knowledge of the dimension for narrow-sense BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ only covers designed distances $2\leq\delta\leq\frac{q^{\lceil(m+1)/2\rceil}-1}{\lambda}+1$ and some specific cases. Our results significantly extend this range, as evidenced by the inequality $\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}\geq\frac{q^{\lfloor(m+1)/2\rfloor}-1}{\lambda}\cdot q^{\lceil(m-4)/6\rceil}$. This implies that the range of $\delta$ for which we provide the dimension of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ is substantially larger than previously established. Additionally, we extend these results to certain non-narrow-sense BCH codes of the same length.
This paper is organized as follows. In Sections II and III, we provide essential preliminaries and review our previous results concerning $q$-cyclotomic cosets modulo $q^{m}-1$. Section IV presents several auxiliary lemmas that will be employed in subsequent sections. Based on the theoretical foundations established in Sections II through IV, we then determine the dimension of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$ in Section V and the Bose distance of $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}$ in Section VI, both by providing explicit formulas. Utilizing these formulas, we present some examples of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ and compare them with the tables of the best known linear codes maintained by Markus Grassl at http://www.codetables.de, which is called Database later in this paper. Furthermore, as an illustration of our main results, Section VII applies these formulas to compute the dimension and Bose distance of BCH codes $\mathcal{C}_{(q,(q^{m}-1)/\lambda,\delta)}$ specifically when $\lambda\delta$ takes the form $aq^{h+k}+b$, where $h=\lfloor m/2\rfloor$, and $k,a,b$ are integers such that $m-2h\leq k\leq\lfloor(2m-1)/3\rfloor-h$, $1\leq a\leq q-1$, $\lambda\leq b\leq q^{m-h-k}$ and $q\nmid b$. Following this, Section VIII extends our analysis to determine the dimension and Bose distance of certain non-narrow-sense BCH codes. Moreover, we identify some non-narrow-sense BCH codes that possess optimal parameters. Finally, Section IX concludes the paper.
@@ II Preliminaries
where $\ell_{a}$ is the smallest positive integer such that $aq^{\ell_{a}}\equiv a\pmod{n}$. It is clear that $|C_{n}(a)|=\ell_{a}$, where $|\cdot|$ denotes the size of a set. Moreover, it is well known that $|C_{n}(a)|=\ell_{a}$ is a divisor of $m=\mathrm{ord}_{n}(q)$. The smallest integer in $C_{n}(a)$ is called the coset leader of $C_{n}(a)$. For convenience, we occasionally refer to it simply as a coset leader modulo $n$. Let $\ell$ be a positive integer, and let $c,d\in[0,n-1]$ be two integers such that $c\leq d$. We define
@@
In particular, we simply denote $\mathcal{L}^{n}_{\ell}(0,n-1)$ by $\mathcal{L}^{n}_{\ell}$.
It is widely known that the minimal polynomial $m_{a}(x)$ of $\beta^{a}$ over $\mathbb{F}_{q}$ is given by
$m_{a}(x)=\prod\limits_{i\in C_{n}(a)}(x-\beta^{i});$
see, for example, [17, Theorem 4.1.1]. Hence, the generator polynomial $g(x)$ of the BCH code $\mathcal{C}_{(q,n,\delta,b)}$ can be expressed as
$g(x)=\prod\limits_{i\in\mathcal{G}}(x-\beta^{i}),\quad\mathcal{G}=\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a).$
Consequently, the dimension of $\mathcal{C}_{(q,n,\delta,b)}$ is
$\mathrm{dim}(\mathcal{C}_{(q,n,\delta,b)})=n-\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)\right|,$ (2)
and the Bose distance $d_{B}$ of $\mathcal{C}_{(q,n,\delta,b)}$ is exactly the largest integer such that
$\bigcup\limits_{a=b}^{b+\delta-2}C_{n}(a)=\bigcup\limits_{a=b}^{b+d_{B}-2}C_{n}(a).$ (3)
In particular, the dimension of narrow-sense BCH codes $\mathcal{C}_{(q,n,\delta)}$ can be given by
@@
Thus, determining the coset leaders modulo $n$ within $1,\delta-1]$ and the sizes of corresponding $q$-cyclotomic cosets allows us to compute the dimension of $\mathcal{C}_{(q,n,\delta)}$. Additionally, for $\delta^{\prime}>\delta$, one can easily verify that
$\bigcup_{a=1}^{\delta-1}C_{n}(a)=\bigcup_{a=1}^{\delta^{\prime}-1}C_{n}(a)$
if and only if none of the integers from $\delta$ to $\delta^{\prime}-1$ is a coset leader. Therefore, the Bose distance of $\mathcal{C}_{(q,n,\delta)}$ equals the smallest coset leader that is not less than $\delta$.
We have a useful observation regarding $q$-cyclotomic cosets, as presented in the following lemma. This result originated from the proof of [41, Lemma 6]. For completeness, we include a proof below.
@@
Noticing that $a\bmod n=a-n\cdot\lfloor\displaystyle\frac{a}{n}\rfloor$ and $\lambda a\bmod\lambda n=\lambda a-\lambda n\cdot\lfloor\displaystyle\frac{\lambda a}{\lambda n}\rfloor$, we can assert that
$\lambda\cdot(a\bmod n)=\lambda a-\lambda n\cdot\lfloor\displaystyle\frac{a}{n}\rfloor=\lambda a-\lambda n\cdot\lfloor\displaystyle\frac{\lambda a}{\lambda n}\rfloor=\lambda a\bmod\lambda n.$
@@
By definition, this holds if and only if $\lambda a$ is the coset leader of $C_{\lambda n}(\lambda a)$. Therefore, the first statement follows.
@@
Let $\mathbb{Z}$ be the set of all integers. For each integer $\lambda$, we denote by $\mathcal{D}_{\lambda}$ the set of all the integers that are divisible by $\lambda$, that is,
@@
Then one can directly derive the following corollary from Lemma 1.